Crew 245 GreenHab Report April 16th

Crew 245 GreenHab Report 04-16-2021

GreenHab Officer: none, report written by Crew Scientist Olivia Ettlin

Environmental control: heater is kept on all day at 65 degrees.

Average temperatures: heat kept to 65 degrees, daily temperatures higher due to sunlight.

Hours of supplemental light: 0 hours

Daily water usage for crops: none

Daily water usage for research and/or other purposes: add 500 mL to hydroponics garden daily, ½ gallon added to regolith feasibility experiment

Water in Blue Tank _____ gallons: 0 gallons

Time(s) of watering for crops: watering regolith experiment in the morning at approx. 10am and adding water to hydroponics garden in the evening at approx. 7pm.

Changes to crops: no crops to maintain

Narrative: The soil in the regolith feasibility experiment is slightly drier than expected. Crew Botanist is monitoring seeds carefully to determine whether it’s too hot in the GreenHab during the day. It’s also possible that more water will need to be added to maintain desired moisture.

The Crew Scientist sealed the leaks in the hydroponics garden and is happy to report that the only water lost by this experiment will be due to evaporation. The tomato cuttings that were wilting yesterday were determined to be completely unsalvageable. Today, we re-started the tomato plant aspect of the hydroponics experiment using rooted tomato cuttings from the same plants. So the growth of these plants will be a day behind the green onions. The Crew Scientist moved the hydroponic garden to the ground floor of the main Hab hopefully to remedy the effects of the extreme heat in the GreenHab.

Harvest: none

Support/supplies needed: none at this time.

-Olivia Ettlin
Red Planet People – MDRS Crew 245 "Team Patamars"

Operations Report April 16th

Crew 245 Operations Report 16-04-2021

SOL: 6

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 6.5

From what time last night: 10:30 pm, SOC 53%

To what time this morning: 5:00 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 69% (Before generator is run at night),

Diesel Reading – 50%

Station Propane Reading – 73%

Water (loft tank): 25 gallons

Water Meter: 150396.0 units

Water (static tank): 300 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 202.1

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.1

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.5 gallon (estimated)

# Hours the ATVs were used today: 0.75 (but 0.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew did not consume allotted bandwidth last night so there was an additional bandwidth bonus available today.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 used on EVA today and functioned nominally.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment. Crew Scientist moved hydroponics experiment from GreenHab to lower deck of Hab, as intense heat and sunlight in GreenHab was causing cuttings to dry up.

Summary of ScienceDome operations: Used by HSO for mini-botany experiment (Space Farmer Dream Kit), with assistance from the Crew Botanist.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

Shravan Hariharan, Crew Engineer

Sol Summary Report

Crew 245 Sol Summary Report 16-04-2021

Sol: 6

Summary Title:

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 08:30-09:30: Jamming out as a crew to Hamilton and Disney music while cooking and eating breakfast, then doing debriefing. Biscuits and gravy for breakfast!

· 09:30-10:30: Preparing gopro and science mission equipment sterilization.

· 10:30: Finish 5 minutes in airlock and EVA Crew #1 out onto Mars.

· 10:30-12:00: Crew 1 out on EVA.

· ~12:00: Crew 1 recalled from EVA.

· 12:30-13:00: Crew spoke with Outpost regarding Map usage over instruments.

· 13:00-14:30: Vegetable pot pie!

· 14:30-17:00: Misc. science.

· 17:00-6:30: Learned how to make challah!

Look Ahead Plan: Tomorrow is EVA #6 and #7, where the crew will practice map usage.

Anomalies in work: early return due to over-reliance on GPS over maps.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #5, with EVA crew #1. Report incoming.

Reports to be filed: Operations Report, EVA Report, EVA Plan

Support Requested: n/a

Copyright © The Mars Society. All rights reserved. | Main Site